# swvicad.bib

### v1

@@ 3.2 Space B: Diving into Spin Waves
The artistic goal of the exhibit is approached by making spin waves audibly and visually perceivable for a human. Under normal circumstances this is impossible due to the scale and speed of the phenomenon, the exhibit features Space B, a special space allowing visitors to overcome limitations of human perception and experience what spin waves could look and sound like, see Fig. 6 Space B. The room has no lights except for the minimum required for the visitors’ safety. One of the walls in the space is occupied by a large screen [Fig. 1 (a)] where visual representations of five different microstrip structures are projected (see Section 2.2). Loudspeakers [Fig. 1 (c)] are placed behind the screen to reinforce the sensation of the sound coming from the strips. About 3 meters in front of the screen five foot pedals, one for each projected microstrip, are set up [Fig. 1 (b)]. By pressing a pedal visitors are able to start the simulation of spin-wave excitation in the respective strip. As a result, the strip visualization starts pulsating following the simulated patterns of magnetization (see Sections 2.1 and 2.2), while the audio representation of the process is played in parallel (see Section 2.3). A video recording of the exhibit showing the path through spaces A (without a brief introduction by a guide) and B including an example of interaction with the pedals is available [^6] https://youtu.be/DVGgFnQUK1g.
@@
The audio-visual sequence of each strip is unique: gradually evolving from the initial state (no spin-waves excited) to a distinct stable state (the spin-wave pattern is stable). The screen recordings of the sequences corresponding to each strip are available online: vase [^7] https://youtu.be/a8Ruy1NVoG0, ellipse [^8] https://youtu.be/bt7djw2yBnA, pyramid [^9] https://youtu.be/CtKyRkgrZQE, rectangular strip [^10] https://youtu.be/tpxx9haPGOE, wave [^11] https://youtu.be/ucMj71vy8bg. The audio-visual display with its loud sound and pulsating imagery turns a humble micro-scale element, one whose countless barely noticeable counterparts serve as building blocks for modern smart devices, into a force to be reckoned with. The contrast between seeing a microstrip being shorter than the width of a human hair and experiencing it as a large object housing complex processes and producing powerful unusual sounds is meant to make the exhibit memorable to the visitors.
@@
Visitors, usually in groups of two or three people, are free to explore the spaces in any order: start with an introduction and then proceed to the display or first interact with the exhibit and then ask questions and get more context or skip any of the parts. The audio-visual display supports simultaneous activation of any number of strips allowing a small group of visitors to interact with the exhibit together and experiment with combining sounds of different shapes. Screen recordings of two strips activated simultaneously are available online: ellipse (panned more to the left channel compared to the exhibit) + wave [^12] https://youtu.be/tTI19KiyDo0, ellipse (panned more to the left channel compared to the exhibit) + rectangular strip [^13] https://youtu.be/JPfCqN6OlNY. In considerations of hygiene and visitor safety foot pedals are selected as a means of interaction with the display. Additionally, the microscope is equipped with a digital camera transmitting the image to a monitor allowing visitors to see the size comparison between the microstrip and a human hair without the need to touch the microscope.
@@ 4 Reception and Discussion
The installation, as described in Section 3, including both spaces and interactive guidance provided by the authors of this paper, was presented at an annual international digital art exhibition Ars Electronica Festival in 2022 with the topic of the year ”Welcome to Planet B”. There Spin-Wave Voices introduced visitors to spin waves as a potential way to make computational devices more energy efficient and thus contribute to solving the global energy crisis.
Additionally, a simplified version of the developed prototype was demonstrated at two research groups, one of which specializes in micromagnetic simulations, and another in research, which also includes measurements of complex magnetization configurations within magnetic materials. The goal was to gather feedback from the specialists in the field.
@@ 4.1 Ars Electronica Festival
The festival spanned over 5 days and attracted 71,000 visitors at 11 locations situated in Linz, Austria [^14] https://ars.electronica.art/mediaservice/en/2022/09/12/71000-visits-to-planet-b/. Spin-Wave Voices was placed in a glass cubicle situated in a hall with other exhibits at the Johannes Kepler University Linz campus. In order to create dark space inside and provide sound insulation the cubicle was covered with thick dark stage curtains around its perimeter from the inside. The same kind of curtain was used to separate the Space A with the microscope and the Space B with the audio-visual display the way that the Space B was not visible from the installation entrance (see Fig. 6). One of the reasons for the separation curtains placement was to prevent additional light in the Space B, another reason was to create a maze-like feeling of the exhibit. The entrance to the cubicle was always open in order to invite visitors in and provide additional light to the room with the microscope. While giving an interactive introduction to the exhibit to all willing participants, the authors did their best to observe their reactions to the installation and collect verbal feedback. The main observations are presented below.
@@
  • Many visitors were willing to first find out more about the exhibit and its meaning before interacting with the audio-visual display. Apparently, the variety of objects present in the Space A helped engage different kinds of participants. While the microscope was appreciated by most of the visitors, the cleanroom suits attracted the attention of a younger audience. At the same time, the audience members with scientific background appreciated the posters providing more detail on the ongoing research of spin waves. The introduction part ended up being key to the exhibit both to convey the main messages of the exhibit but also to engage the visitors with the narrative.
@@
From the observations listed above the following conclusion can be drawn: the first three objectives listed in Section 1.4 were reached, namely the public awareness of the need for energy-efficient computational technologies, e.g. spin waves was raised; the visitors got an opportunity to dive deeper into nano-scale ultrafast processes akin to ones behind modern and future computational technologies; and an accessible immersive audio-visual experience was provided, which illustrates the basic concept of spin waves in computing. The fourth objective to explore the potential of sonification as a tool for analysis of spin-wave data is discussed in Section 4.2).
Additionally, from the listed observations the following improvements might help to get an even better reach in the future: (1) an audio-based exhibit needs ways to attract the public when the setup is not in action, (2) artistic and science communication objectives can clash, for example when visitors do not engage with the message of the exhibit because of its looks or sound being alien to them, (3) providing explicit interactive introduction can increase visitor engagement with the exhibit, (4) special care needs to be taken to prevent potential misinterpretation of a science communication exhibit.
@@ 4.2 Demonstration to Specialists in the Field
A simplified version of Spin-Wave Voices (pedals were replaced by a keyboard) of the developed prototype was presented in two research groups, one of which specializes in micromagnetic simulations, and another in research, which also includes measurements of complex magnetization configurations within magnetic materials. The idea was suggested as a potential augmentation of visual analysis of complex spin-wave data. The feedback from the specialists is summarized in the following list:
@@
Therefore, the fourth objective listed in Section 1.4, namely to explore the potential of sonification as a tool for analysis of spin-wave data, was reached as well. If developed further, the approach can find its application in the magnonics research field and further, in the research of the complex magnetic phenomena. With the little improvement, e.g. allowing to easily switch the sonified data and adjust the parameters of the visualization and, hence, AD, even the current approach could already be used as a tool for spin-wave data analysis.
@@ 5 Conclusion and Outlook
This paper introduces an approach to constructing an audio-visual representation of complex spin-wave data. The work at hand showcases the use of the approach as a part of a science communication art exhibit and outlines steps to be taken to develop it into an efficient tool for the analysis of spin-wave behavior, allowing researchers to rely both on their vision and hearing for discovering new patterns in the data. Spin-Wave Voices featuring the presented approach was successfully demonstrated at the annual international digital art exhibition Ars Electronica Festival. The paper reports details of the setup as well as observations on the visitor reception of the exhibit. Additionally, results of the presentation to the specialists in the magnetic phenomena research are discussed.
One direction of future work is to develop the approach into an independent tool for spin-wave data analysis. This would require unifying the code/software base, e.g. turning the display into a single Python application that would allow for easier deployment and integration with other tools. A study would be required to determine the most useful modes of sonification, each aimed at its own analysis task. Finally, the tool will need to provide flexibility in terms of sonification: manual path selection, cropping of the data, and area aggregation parameters.
The other direction would be to make the exhibit based on the presented audio-visual display more engaging and informative, which is important for communicating spin-wave research results not only to the general public, schools, the media, and policy makers, but also to the scientific community itself. Boosting visitor engagement could be achieved by introducing an attractive stand-by mode and giving visitors more control over the installation, e.g. choosing base frequencies and other sonification parameters for each strip. In turn, addressing visualization ambiguity and demonstrating spin waves in comparison to modern technology, e.g. showcasing differences in produced heat through engaging thermoreception of the visitors, could make the exhibit more informative.

### v2

@@ 3.2 Space B: Diving into Spin Waves
The artistic goal of the exhibit is approached by making spin waves audibly and visually perceivable for a human, which is impossible under normal circumstances due to the scale and speed of the phenomenon. Therefore, the exhibit features Space B, a special space allowing visitors to overcome limitations of human perception and experience what spin waves could look and sound like, see Space B in Fig. 6. The room has no lights except for the minimum required for the visitors’ safety. One of the walls in the space is occupied by a large screen [Fig. 1 (a)] where visual representations of five different microstrip structures are projected (see Section 2.2). Loudspeakers [Fig. 1 (c)] are placed behind the screen to reinforce the sensation of the sound coming from the strips. About 3 meters in front of the screen five foot pedals, one for each projected microstrip, are set up [Fig. 1 (b)]. By pressing a pedal visitors are able to start the simulation of spin-wave excitation in the respective strip. As a result, the strip visualization starts pulsating following the simulated patterns of magnetization (see Sections 2.1 and 2.2), while the audio representation of the process is played in parallel (see Section 2.3). A video recording of the exhibit showing the path through spaces A (without a brief introduction by a guide) and B including an example of interaction with the pedals is available [^6] https://youtu.be/DVGgFnQUK1g.
@@
The audio-visual sequence of each strip is unique: gradually evolving from the initial state (no spin waves excited) to a distinct stable state (the spin-wave pattern is stable). The screen recordings of the sequences corresponding to each strip are available online: vase [^7] https://youtu.be/a8Ruy1NVoG0, ellipse [^8] https://youtu.be/bt7djw2yBnA, pyramid [^9] https://youtu.be/CtKyRkgrZQE, rectangular strip [^10] https://youtu.be/tpxx9haPGOE, wave [^11] https://youtu.be/ucMj71vy8bg. The audio-visual display with its loud sound and pulsating imagery turns a humble micro-scale element, one whose countless barely noticeable counterparts serve as building blocks for modern smart devices, into a force to be reckoned with. The contrast between seeing a microstrip being hundred times smaller than the width of a human hair and experiencing it as a large object housing complex processes and producing powerful unusual sounds is meant to make the exhibit memorable to the visitors.
@@
Visitors, usually in groups of two or three people, are free to explore the spaces in any order: start with an introduction and then proceed to the display or first interact with the exhibit and then ask questions and get more context or skip any of the parts. The audio-visual display supports simultaneous activation of any number of strips allowing a small group of visitors to interact with the exhibit together and experiment with combining sounds of different shapes. Screen recordings of two strips activated simultaneously are available online: ellipse (panned more to the left channel compared to the exhibit) + wave [^12] https://youtu.be/tTI19KiyDo0, ellipse (panned more to the left channel compared to the exhibit) + rectangular strip (panned more to the right channel compared to the exhibit)  [^13] https://youtu.be/JPfCqN6OlNY. In considerations of hygiene and visitor safety foot pedals are selected as a means of interaction with the display. Additionally, the microscope is equipped with a digital camera transmitting the image to a monitor allowing visitors to see the size comparison between the microstrip and a human hair without the need to touch the microscope.
@@ 4 Reception and Discussion
The installation, as described in Section 3, including both spaces and interactive guidance provided by the authors of this paper, was presented at an annual international digital art exhibition Ars Electronica Festival in 2022 with the topic of the year “Welcome to Planet B”. There Spin-Wave Voices introduced visitors to spin waves as a potential way to make computational devices more energy efficient and, thus, contribute to solving the global energy crisis. Additionally, a simplified version of the developed prototype was demonstrated at two research groups, one of which specializes in micromagnetic simulations, and another in research, which also includes measurements of complex magnetization configurations within magnetic materials, and at the INTERMAG 2024 as a contributed talk. The goal was to gather feedback from the specialists in the field.
@@ 4.1 Ars Electronica Festival
The festival spanned over 5 days and attracted 71,000 visitors at 11 locations situated in Linz, Austria [^14] https://ars.electronica.art/mediaservice/en/2022/09/12/71000-visits-to-planet-b/. Spin-Wave Voices was placed in a glass cubicle situated in a hall with other exhibits at the Johannes Kepler University Linz campus. In order to create dark space inside and provide sound insulation the cubicle was covered with thick dark stage curtains around its perimeter from the inside. The same kind of curtain was used to separate the Space A with the microscope and the Space B with the audio-visual display the way that the Space B was not visible from the installation entrance (see Fig. 6). One of the reasons for the separation curtains placement was to prevent additional light in the Space B, another reason was to create a maze-like feeling of the exhibit. The entrance to the cubicle was always open in order to invite visitors in and provide additional light to the room with the microscope [^15] https://youtu.be/DVGgFnQUK1g.
While giving an interactive introduction to the exhibit to all willing participants, the authors did their best to observe their reactions to the installation and collect verbal feedback. As no means of formal evaluation was available at the moment of the exhibition, below we list observations all five authors found relevant.
@@
  • Many visitors were willing to first find out more about the exhibit and its meaning before interacting with the audio-visual display. Apparently, the variety of objects present in the Space A, the microscope, the cleanroom suits, and the scientific poster with more detail on the ongoing research of spin waves, helped engage different kinds of participants. The introduction part ended up being key to the exhibit both to convey the main messages of the exhibit but also to engage the visitors with the narrative.
@@
From the observations listed above the following conclusion can be drawn: the first three objectives listed in Section 1.4 were reached, namely the public awareness of the need for energy-efficient computational technologies, e.g. spin waves was raised (see objective (1) in Section 1.4); the visitors got an opportunity to dive deeper into nano-scale ultrafast processes akin to ones behind modern and future computational technologies (see objective (2) in Section 1.4); and an accessible immersive audio-visual experience was provided, which illustrates the basic concept of spin waves in computing (see objective (3) in Section 1.4). The fourth objective to explore the potential of sonification as a tool for analysis of spin-wave data is discussed in Section 4.2).
Given that the conclusions stem from informal discussions with the visitors, in the future work it is planned to conduct formal evaluation of the setup to quantify participant initial interest, engagement and learning effect (e.g. through monitoring of objective parameters, such as time spent interacting with each part of the exhibit, as well as via questionnaires). Nevertheless, the listed observations have already shown that the following improvements might help to get an even better reach in the future: (1) an audio-based exhibit needs ways to attract the public when the setup is not in action, (2) artistic and science communication objectives can clash, for example when visitors do not engage with the message of the exhibit because of its looks or sound being alien to them, (3) providing explicit interactive introduction can increase visitor engagement with the exhibit, (4) special care needs to be taken to prevent potential misinterpretation of a science communication exhibit.
@@ 4.2 Demonstration to Specialists in the Field
The idea of Spin-Wave Voices with short demonstration was presented in two research groups, one of which specializes in micromagnetic simulations, and another in research, which also includes measurements of complex magnetization configurations within magnetic materials. Additionally, similar presentation was done at the INTERMAG 2024 as a contributed talk. The idea of the project was suggested as a potential augmentation of visual analysis of complex spin-wave data. The feedback from the specialists is summarized in the following list:
@@
Therefore, the objective (4) in Section 1.4, namely to explore the potential of sonification as a tool for analysis of spin-wave data, was reached as well. If developed further, the approach can find its application in the magnonics research field and further, in the research of the complex magnetic phenomena. With the little improvement, e.g. allowing to easily switch the sonified data and adjust the parameters of the visualization and, hence, AD, even the current approach could already be used as a tool for spin-wave data analysis.
@@ 5 Conclusion and Outlook
This paper introduces an approach to constructing an audio-visual representation of complex spin-wave data. The work at hand showcases the use of the approach as a part of a science communication art exhibit and outlines steps to be taken to develop it into an efficient tool for the analysis of spin-wave behavior, allowing researchers to rely both on their vision and hearing for discovering new patterns in the data. Spin-Wave Voices featuring the presented approach was successfully demonstrated at the annual international digital art exhibition Ars Electronica Festival 2022 “Welcome to Planet B”, where it has shown potential to communicate recent advances in spin-wave research, which aims for more energy-efficient computational devices, to general public (see objective (1) in Section 1.4), served as an engaging illustration to spin waves as a process beyond human perception (due to its scale and speed) (see objective (2) in Section 1.4) and provided a captivating audio-visual experience to the visitors (see objective (3) in Section 1.4). The paper reports details of the setup as well as observations on the visitor reception of the exhibit. Additionally, results of the presentation to the specialists in the magnetic phenomena research are discussed. This served as an initial proof of concept encouraging further investigation of ways to leverage sonification for spin-wave data analysis (see objective (4) in Section 1.4).
Given that the conclusions presented in this paper stem from informal discussions with the visitors of the digital art festival and the physicists, in the future work it is planned to conduct formal evaluation of the setup as a science-communication tool with the main objective to quantify its attractiveness and effectiveness. The results of the evaluation will be the guiding factor to improving the exhibit: making sure it conveys intended messages, mitigating interaction bottlenecks and boosting accessibility. The next steps in the development of the exhibit based on the presented audio-visual display include making it more engaging and informative and expanding its scope from general public to schools, the media, policy makers, as well as to the scientific community. Boosting visitors engagement could be achieved by introducing an attractive stand-by mode and giving visitors more control over the installation, e.g. choosing base frequencies and other sonification parameters for each strip. In turn, addressing visualization ambiguity and demonstrating spin waves in comparison to modern technology, e.g. showcasing differences in produced heat through engaging thermoreception of the visitors, could make the exhibit more informative. Another direction of future work is to develop the approach into an independent tool for spin-wave data analysis. This would require unifying the code/software base, e.g. turning the display into a single Python application that would allow for easier deployment and integration with other tools. A study would be required to determine the most useful modes of sonification, each aimed at its own analysis task. Finally, the tool will need to provide flexibility in terms of sonification: manual path selection, cropping of the data, and area aggregation parameters.